OpenStudy (anonymous):

solve the inequality and graph the solution on the real line. 0≤x+3<5

OpenStudy (helder_edwin):

add -3 to this

OpenStudy (anonymous):

to what?

OpenStudy (helder_edwin):

your inequality \[ \large 0\leq x+3<5 \]

OpenStudy (anonymous):

yes where do i put it?

OpenStudy (anonymous):

like I'm not sure where to start

OpenStudy (helder_edwin):

what?

OpenStudy (helder_edwin):

\[ \large 0\leq x+3<5\qquad//-3 \] \[ \large 0+(-3)\leq x+3+(-3)<5+(-3) \]

OpenStudy (anonymous):

So its -3<x<2 ????

OpenStudy (helder_edwin):

no \[ \large -3\leq x <2 \]

OpenStudy (anonymous):

yeah thats what i had..are u sure?

OpenStudy (helder_edwin):

yes

OpenStudy (anonymous):

how would i do the second part of prob?

OpenStudy (helder_edwin):